# SDP bounds on quantum codes

### v1

@@ 7. Quantum Delsarte bound
For classical codes, the symmetrization over $\operatorname{Aut}(q,n)$ of the semidefinite program SDP1 for the Lovász theta number [Eq. (3.5)] along with the non-negativity of the matrix entries as extra constraint leads to the Delsarte bound [4]. The matrix used for the Lovász theta bound is defined as $R=\chi\chi^{T}$, where $\chi$ is the indexing vector of a classical code. Our aim is to recover a similar result for the Lovász number for self-dual quantum codes. For simplicity in the presentation, we here use the semidefinite program SDP2 [see Eq. (3.5)]. A key difference to the classical case is that the entries of the semidefinite variable from the quantum Lovász theta bound [see Corollary 9] can be negative. Due to this fact, we require a different extra constraint to the quantum Lovász theta bound to recover the quantum Delsarte bound [see Eq. (3.4)] after the symmetrization.
Suppose $\Gamma$ arises from a state. Our aim is to show that the $A_{j}(\rho)$ enumerators emerge as elements from $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ expressed in the Bose-Mesner algebra basis $\{N_{k}\}_{k=0}^{n}$ [see Eq. (90)]. Diagonalizing $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$, one then obtain by means of the quantum MacWilliams identity, the dual enumerator $B_{j}(\rho)$. As in Eq. (88), the matrix $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ is part of the Bose-Mesner algebra and can be expanded in terms of the basis $\{M^{t,p}_{i,j}\}_{(i,j,t,p)\in\mathcal{I}(4,n)}$ of the Terwilliger algebra,
@@
Recall that $M^{t,p}_{i,j}$ is an orthogonal basis, that is $\left\langle M_{i,j}^{t,p},M_{i,j}^{t,p}\right\rangle=\gamma_{i,j}^{t,p}$ with $\gamma_{i,j}^{t,p}$ given by Eq. (77). Since all $M_{i,j}^{t,p}$ with $k=i+j-t-p$ have the same coefficient $y_{k}$, it is enough to compute the inner product with a representative, e.g., $M_{k,0}^{0,0}$,
[…]
Deriving ${\langle\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu,M^{0,0}_{k,0}\rangle}$ is thus enough to determine $y_{k}$. Following the procedure of Eq. (6.3) to (83) and using the definition of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ in Eq. (96), one obtains
@@
We now show that $y_{k}$ is proportional to the $A_{k}(\rho)$ enumerator. Recall the action of $\sigma_{v}$ on $\Gamma$ in Eq. (6.5). To this end, let us expand for a quantum code $\varrho=\Pi/K$ the last sum in Eq. (7),
@@
Here we have used that $\varrho^{2}=\varrho/K$ and $\operatorname{tr}((1,2)\varrho^{\otimes 2})=\varrho^{2}$ where the swap operator is $(1,2)=2^{-n}\sum_{E_{v}\in\mathcal{E}_{n}}E_{v}\otimes E^{\dagger}_{v}$. Note that Eq. (7) can also be seen from combining the constraints Eq. (5.2) and Eq. (62) in the main semidefinite programming relaxation.
@@
We now diagonalize $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$. Ref [13] shows that the eigenvalues from $N_{k}$ are the Krawtchouk polynomials $K_{k}(i)$ [see Eq. (16)]. Since $\{N_{k}\}^{n}_{k=0}$ is a commuting basis it can be simultaneous diagonalized, thus the eigenvalues of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ are $\sum^{n}_{k=0}y_{k}K_{k}(i)$. The quaternary Krawtchouk polynomials satisfy the following relation [44, Theorem 17],
(102) $3^{i}\binom{n}{i}K_{k}(i)=3^{k}\binom{n}{k}K_{i}(k)\,.$
With Eq. (77), one can express this as $\gamma^{0,0}_{i,0}K_{k}(i)=\gamma^{0,0}_{k,0}K_{i}(k)$. We then develop the eigenvalues of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ as follows
$\sum^{n}_{k=0}y_{k}K_{k}(i)$ $=\frac{1}{\gamma^{0,0}_{i,0}}\sum^{n}_{k=0}y_{k}\gamma^{0,0}_{k,0}K_{i}(k)$
(103) $=(2^{n}K\gamma^{0,0}_{k,0})^{-1}\sum^{n}_{k=0}K_{i}(k)A_{k}(\varrho)\,,$
@@
However, we want to point out that Eq. (7) assumes that $\Gamma$ has been constructed from a state $\varrho=\Pi/K$. Thus this is an extra constraint that has to be imposed on the Lovász bound in order to obtain the quantum Delsarte bound:
@@
(104) $\sum_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}=2^{n}\Gamma_{x0}\,,$
@@
(107) $(2^{n}\gamma^{0,0}_{j,0})^{-1}\sum^{n}_{j=0}K_{i}(j)A_{j}\,.$
Therefore, $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\succeq 0$ if and only if $\sum^{n}_{j=0}K_{i}(j)A_{j}\geq 0$ for all $0\leq i\leq n$. The combination with (1) - (4) above yields the quantum Delsarte bound for self-dual codes [Eq. (3.4)],
@@
(108) $\sum^{n}_{i=0}K_{j}(i)a_{i}\geq 0\quad\text{for}\quad 0\leq j\leq n\,.$
@@
Naturally, this bound can be made stronger by formulating it as a feasibility program that includes the constraint $\sum_{j=0}^{n}a_{j}=2^{n}$. We use this for excluding the $(\!(7,1,4)\!)_{2}$ code in Section 9.1.
@@ 8. Symmetry-reduced semidefinite programming bounds
We now aim to symmetry-reduce $\Gamma$ while retaining as many constraints from the original program [see Eq. (5.7)] as possible. This is achieved by averaging $\Gamma$ over $\operatorname{Aut}_{0}(4,n)$, which corresponds to Eq. (95) with $Q=\{0^{n}\}$. We therefore define
@@
In analogy to Eq. (87), we define $\tilde{\Gamma}^{c}$ as
(110) $\tilde{\Gamma}^{c}=\frac{1}{4^{n}-2^{n}/K}\left(4^{n}\bar{\Gamma}-\frac{2^{n}}{K}\tilde{\Gamma}\right)\,.$
where $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ is stated in Eq. (96). Note that we choose $2^{n}/K$ in the normalization of $\tilde{\Gamma}$, which in case of stabilizer codes this corresponds to the number of stabilizers. In general however, the matrix $\tilde{\Gamma}^{c}$ does not necessarily come from averaging $\Gamma$ over the codewords of a code, i.e., $\tilde{\Gamma}^{c}\neq\frac{1}{4^{n}-\left\lvert Q\right\rvert}\sum_{v\notin Q}\Gamma^{v}$ for some $Q\subseteq{\boldsymbol{4}}^{n}$. It is thus not immediate that $\tilde{\Gamma}^{c}$ is positive semidefinite. We now show that $\tilde{\Gamma}^{c}\succeq 0$ for all quantum codes.
@@
is positive semidefinite. To see this, write $L=M\circ G$ where $\circ$ is the Schur product, that is $L_{xy}=M_{xy}G_{xy}$, and where
@@
As it has the form of a moment matrix it is clear that $G\succeq 0$. Let us show that also $M\succeq 0$. Define two matrices $M^{\pm}$ with entries
(113) $M^{\pm}_{xy}=2\operatorname{tr}\Big{(}R_{x}R^{\dagger}_{y}\big{(}[\operatorname{id}\pm(1,2)]\varrho^{\otimes 2}\big{)}\Big{)}\quad\text{with}\quad R_{x}:=\frac{1}{2}[\operatorname{id}\mp(1,2)](E_{x}\otimes\operatorname{id})\,.$
Now we use the fact that $\frac{1}{2}[\operatorname{id}\pm(1,2)]$ is a projector and that $(1,2)A\otimes B=B\otimes A(1,2)$. Then $M^{\pm}\succeq 0$ when $\varrho\succeq 0$, because 666 This is inspired by Ref. [59, Lemma 7.2], where for any operator $A$, the inequality $(A\otimes A^{\dagger}(1,2))^{T_{2}}\succeq 0$ is derived. Here $T_{2}$ is the partial transpose in the second subsystem.
(114) $\frac{1}{2}[\operatorname{id}\pm(1,2)]\varrho^{\otimes 2}=\frac{1}{2}[\operatorname{id}\pm(1,2)]\,\,\varrho^{\otimes 2}\,\,\frac{1}{2}[\operatorname{id}\pm(1,2)]\succeq 0\,.$
Expand the matrix $M^{\pm}$ to obtain
$M^{\pm}_{xy}$ $=\operatorname{tr}\Big{(}(E_{x}\otimes\operatorname{id})(1\mp(1,2))(E^{\dagger}_{y}\otimes\operatorname{id})\big{(}[\operatorname{id}\pm(1,2)]\varrho^{\otimes 2}\big{)}\Big{)}$
$=\operatorname{tr}([E_{x}E^{\dagger}_{y}\otimes\operatorname{id}]\varrho^{\otimes 2})\mp\operatorname{tr}([E_{x}\otimes E^{\dagger}_{y}](1,2)\varrho^{\otimes 2})$
±$\quad\pm\operatorname{tr}([E_{x}E^{\dagger}_{y}\otimes\operatorname{id}](1,2)\varrho^{\otimes 2})-\operatorname{tr}([E_{x}\otimes E^{\dagger}_{y}]\varrho^{\otimes 2})$
(115) $=\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)\operatorname{tr}(\varrho)\mp\operatorname{tr}(E_{x}\varrho E^{\dagger}_{y}\varrho)\pm\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho^{2})-\operatorname{tr}(E_{x}\varrho)\operatorname{tr}(E^{\dagger}_{y}\varrho)\,,$
The average of both matrices reads
(116) $\frac{M^{+}_{xy}+M^{-}_{xy}}{2}=\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)-\operatorname{tr}(E_{x}\varrho)\operatorname{tr}(E^{\dagger}_{y}\varrho)\succeq 0\,,$
which are the entries of $M^{T}$. Therefore, $M\succeq 0$ and also $L=M\circ G\succeq 0$.
@@
(120) $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu_{xy}$ $=\frac{1}{(2^{n}K)\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(S\right)\,,$
@@ 8.1. Positive semidefinite constraints in the Terwilliger algebra
To make use of Proposition 16 in a symmetry-reduced semidefinite program, we need to express these constraints in terms of the Terwilliger algebra.
@@
The matrices $\tilde{\Gamma}$ and $\tilde{\Gamma}^{c}$ can in the Terwilliger algebra be expanded as
@@
The matrix $\tilde{\Gamma}$ is invariant under $\operatorname{Aut}_{0}(q,n)$, and consequentially it can be expanded in the basis $M^{t,p}_{i,j}$ of the Terwilliger Algebra. With the definition from Eq. (83), the parameters $\lambda^{t,p}_{i,j}$ are then given by Eq. (124) and $x^{t,p}_{i,j}=\frac{\lambda^{t,p}_{i,j}}{\gamma^{t,p}_{i,j}}$.
From Eq. (110), we can express $\tilde{\Gamma}^{c}$ in the Terwilliger Algebra basis using Eq. (97),
@@
(126) $y_{k}=(2^{n}K\gamma^{0,0}_{k,0})^{-1}\sum_{\begin{subarray}{c}x\in{\boldsymbol{4}}^{n}\\ |s(x)|=k\end{subarray}}\Gamma_{x0}=\frac{2^{n}}{K}x^{0,0}_{k,0}\,,$

### v2

@@ 7. Quantum Delsarte bound
For classical codes, the symmetrization over $\operatorname{Aut}(q,n)$ of the semidefinite program SDP1 for the Lovász theta number [Eq. (3.5)], along with the non-negativity of the matrix entries as an extra constraint, leads to the Delsarte bound [4]. The matrix used for the Lovász theta bound is defined as $R=\chi\chi^{T}$, where $\chi$ is the indexing vector of a classical code. Our aim is to recover a similar result for the Lovász theta number for self-dual quantum codes. For simplicity in the presentation, we here use the semidefinite program SDP2 [see Eq. (3.5)]. A key difference to the classical case is that the entries of the semidefinite variable from the quantum Lovász theta bound [see Corollary 10] can be negative. Due to this fact, we require a different extra constraint to the quantum Lovász theta bound to recover the quantum Delsarte bound [see Eq. (3.4)] after the symmetrization.
Suppose $\Gamma$ arises from a state. Our aim is to show that the $A_{j}(\rho)$ enumerators emerge as elements from $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ expressed in the Bose-Mesner algebra basis $\{N_{k}\}_{k=0}^{n}$ [see Eq. (92)]. Diagonalizing $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$, one then obtains the dual enumerator $B_{j}(\rho)$ by means of the quantum MacWilliams identity. As in Eq. (90), the matrix $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ is part of the Bose-Mesner algebra and can be expanded in terms of the basis $\{M^{t,p}_{i,j}\}_{(i,j,t,p)\in\mathcal{I}(4,n)}$ of the Terwilliger algebra,
@@
Recall the matrices $M^{t,p}_{i,j}$ form an orthogonal basis, and that $\left\langle M_{i,j}^{t,p},M_{i,j}^{t,p}\right\rangle=\gamma_{i,j}^{t,p}$, where $\gamma_{i,j}^{t,p}$ is given by Eq. (79). Since all $M_{i,j}^{t,p}$ with k=i+jtpk=i+j-t-p have the same coefficient $y_{k}$, it is enough to compute the inner product with a representative, e.g., $M_{k,0}^{0,0}$,
(100) $y_{k}$ $=\frac{1}{\gamma^{0,0}_{k,0}}{\langle\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu,M^{0,0}_{k,0}\rangle}=\frac{1}{\gamma^{t,p}_{i,j}}{\langle\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu,M^{t,p}_{i,j}\rangle}\quad\quad\text{for all}\quad k=i+j-t-p\,.$
To determine $y_{k}$ it is therefore sufficient to compute $\langle\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu,M^{0,0}_{k,0}\rangle$}. Following the procedure of Eq. (6.3) and (85) and using the definition of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ in Eq. (98), one obtains
@@
We now show that $y_{k}$ is proportional to the $A_{k}(\rho)$ enumerator. Recall the action of $\sigma_{v}$ on $\Gamma$ in Eq. (6.5). To this end, let us expand for a quantum code $\varrho=\Pi/K$ the inner sum of Eq. (7),
@@
Here we used that $\varrho^{2}=\varrho/K$ and $\operatorname{tr}\big{(}(1,2)\varrho^{\otimes 2}\big{)}=\operatorname{tr}(\varrho^{2})$ where the swap operator is $(1,2)=2^{-n}\sum_{E_{v}\in\mathcal{E}_{n}}E_{v}\otimes E^{\dagger}_{v}$. Note that Eq. (7) can also be seen from combining the constraints of Eqs. (5.2) and  (63) in the main semidefinite programming relaxation.
@@
We now diagonalize $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$. Ref [13] shows that the eigenvalues of $N_{k}$ can be derived from the Krawtchouk polynomials $K_{k}(i;n)$ [see Eq. (16)]. Since $\{N_{k}\}^{n}_{k=0}$ is a commuting basis it can be simultaneous diagonalized, thus the eigenvalues of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ are $\sum^{n}_{k=0}y_{k}K_{k}(i;n)$. The quaternary Krawtchouk polynomials satisfy the following relation [45, Theorem 17],
(104) $3^{i}\binom{n}{i}K_{k}(i;n)=3^{k}\binom{n}{k}K_{i}(k;n)\,.$
With Eq. (79), one can express this as $\gamma^{0,0}_{i,0}K_{k}(i;n)=\gamma^{0,0}_{k,0}K_{i}(k;n)$. We then develop the eigenvalues of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ as follows
$\sum^{n}_{k=0}y_{k}K_{k}(i;n)$ $=\frac{1}{\gamma^{0,0}_{i,0}}\sum^{n}_{k=0}y_{k}\gamma^{0,0}_{k,0}K_{i}(k;n)$
(105) $=(2^{n}K\gamma^{0,0}_{k,0})^{-1}\sum^{n}_{k=0}K_{i}(k;n)A_{k}(\varrho)\,$,
@@
However, we want to point out that Eq. (7) assumes that $\Gamma$ has been constructed from a state $\varrho=\Pi/K$, for which Eq. (7) applies. We impose this extra constraint on the Lovász bound to obtain the quantum Delsarte bound:
@@
(106) $\sum_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}=2^{n}\Gamma_{x0}\quad\forall x\in{\boldsymbol{4}}^{n}\,$,
@@
(109) $(2^{n}\gamma^{0,0}_{j,0})^{-1}\sum^{n}_{j=0}K_{i}(j;n)A_{j}\,.$
Therefore, $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\succeq 0$ if and only if $\sum^{n}_{j=0}K_{i}(j;n)A_{j}\geq 0$ for all $0\leq i\leq n$. The combination with (1) - (4) above yields the quantum Delsarte bound for self-dual codes [Eq. (3.4)],
@@
(110) $\sum^{n}_{i=0}K_{j}(i;n)a_{i}\geq 0\quad\text{for}\quad 0\leq j\leq n\,.$
@@
Naturally, this bound can be made stronger by formulating it as a feasibility program that includes the constraint $\sum_{j=0}^{n}a_{j}=2^{n}$. We use this for excluding the existence of the $(\!(7,1,4)\!)_{2}$ code in Section 9.1.
@@ 8. Symmetry-reduced semidefinite programming bounds
We now aim to symmetry-reduce $\Gamma$ while retaining as many constraints from the original program [Eq. (5.7)] as possible. This is achieved by averaging $\Gamma$ over $\operatorname{Aut}_{0}(4,n)$, which corresponds to Eq. (97) with $Q=\{0^{n}\}$. We therefore define
@@
In analogy to Eq. (89), define $\tilde{\Gamma}^{c}$ as
[…]
where $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ is stated in Eq. (98). We choose $2^{n}/K$ in the normalization of $\tilde{\Gamma}$, which for stabilizer codes corresponds to the number of stabilizer elements. In general however, the matrix $\tilde{\Gamma}^{c}$ does not necessarily come from averaging $\Gamma$ over the codewords of a code, i.e., $\tilde{\Gamma}^{c}\neq\frac{1}{4^{n}-\left\lvert Q\right\rvert}\sum_{v\notin Q}\Gamma^{v}$ for some $Q\subseteq{\boldsymbol{4}}^{n}$. It is thus not immediate that $\tilde{\Gamma}^{c}$ is positive semidefinite. We now show that $\tilde{\Gamma}^{c}\succeq 0$ for all quantum codes.
@@
is positive semidefinite. To see this, write $L=M\circ G$ where $\circ$ is the Schur product, that is, $L_{xy}=M_{xy}G_{xy}$, and where
@@
As $G$ has the form of a moment matrix it is clear that $G\succeq 0$. Let us show that also $M\succeq 0$. This can be seen by noting that $M$ is a covariance matrix,
$M_{xy}$ $=\big{\langle}\big{(}E_{y}-{\langle E_{y}\rangle}_{\varrho}\big{)}\big{(}E^{\dagger}_{x}-{\langle E^{\dagger}_{x}\rangle}_{\varrho}\big{)}\big{\rangle}_{\varrho}$
(115) $={\langle E_{y}E^{\dagger}_{x}\rangle}_{\varrho}-{\langle E_{y}\rangle}{\langle E^{\dagger}_{x}\rangle}_{\varrho}\,$,
where we used that all observables are Hermitian. Thus $M\succeq 0$ and also $L=M\circ G\succeq 0$.
@@
(119) $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu_{xy}$ $=\frac{1}{2^{n}K\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(S\right)\,$,
@@ 8.1. Positive semidefinite constraints in the Terwilliger algebra
To make use of Proposition 17 in a symmetry-reduced semidefinite program, we need to express the constraints in terms of the Terwilliger algebra.
@@
The matrices $\tilde{\Gamma}$ and $\tilde{\Gamma}^{c}$ expand in the Terwilliger algebra as
@@
The matrix $\tilde{\Gamma}$ is invariant under $\operatorname{Aut}_{0}(q,n)$, and consequentially it can be expanded in the basis elements $M^{t,p}_{i,j}$ of the Terwilliger algebra. With the definition from Eq. (85), the parameters $\lambda^{t,p}_{i,j}$ are then given by Eq. (123) and $x^{t,p}_{i,j}=\frac{\lambda^{t,p}_{i,j}}{\gamma^{t,p}_{i,j}}$.
We can express $\tilde{\Gamma}^{c}$ from Eq. (112) using Eq. (99),
@@
(125) $y_{k}=(2^{n}K\gamma^{0,0}_{k,0})^{-1}\sum_{\begin{subarray}{c}x\in{\boldsymbol{4}}^{n}\\ |s(x)|=k\end{subarray}}\Gamma_{x0}=\frac{2^{n}}{K}x^{0,0}_{k,0}\,.$